Week of 190715

WLCG Operations Call details

  • For remote participation we use the Vidyo system. Instructions can be found here.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 Geneva time until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • General information about the WLCG Service can be accessed from the Operations Portal
  • Whenever a particular topic needs to be discussed at the operations meeting requiring information from sites or experiments, it is highly recommended to announce it by email to wlcg-scod@cernSPAMNOTNOSPAMPLEASE.ch to allow the SCOD to make sure that the relevant parties have the time to collect the required information, or invite the right people at the meeting.

Best practices for scheduled downtimes

Monday

Attendance:

  • local: Alberto (Monitoring), Andreas (Storage), Aritz (Storage), Borja (Chair, Monitoring), Gavin (Computing), Maarten (ALICE), Vincent (Security)
  • remote: Alexander (NL-T1), Darren (RAL), Di (TRIUMF), Sang-Un (KISTI), Ville (NDGF), Wojtek (LHCb)

Experiments round table:

  • ATLAS reports ( raw view) -
    • running with ~ 300k job slots (no CERN-P1 - technical run)
      • more efficient startup of P1 grid resources once they are available
    • pilot 2 migration in progress (using containers if available / working)
    • panda / harvester tuning (exhusted random number generator)
    • user by accident submitting multi-core jobs to single core queue - implemented protection in WMFS

  • ALICE -
    • NTR

  • LHCb reports ( raw view) -
    • Activity:
      • MC, user jobs and staging.
    • Issues:
      • CNAF: All data transfers Failed at INFN-T1; ticket GGUS:142239; fixed

Sites / Services round table:

  • ASGC: NC
  • BNL: NTR
  • CNAF: Issue with 1 VM caused transfer failures at CNAF in principle concerning LHCb. THe machine was moved to another hypervisor. Solved.
  • EGI: NC
  • FNAL: NC
  • IN2P3: NTR
  • JINR: NC
  • KISTI: NTR
  • KIT: NTR
  • NDGF: NTR
  • NL-T1: NTR
  • NRC-KI: NC
  • OSG: NC
  • PIC: NC
  • RAL: NTR
  • TRIUMF: NTR

  • CERN computing services:
    • OTG:0051185 - Hypervisor reboot schedule to mitigate MDS Intel CPU vulnerabilities
      • Will proceed by availability zone, from 18th July. Please watch OTG for updates.
      • Expect CERN CEs to be rebooted on Thursday and Wednesday, OTG:0051291 and OTG:0051294
  • CERN storage services:
    • EOS :
      • OTG0051324 : Due to hypervisor reboot campaign for mds mitigation, some srm and http endpoint will be unavailable
      • EOSALICE : OTG0051349,OTG0051325 : crashes on saturday, EOS-3592 opened to follow the problem.
    • CASTOR: the CASTOR nameserver intervention originally planned for Thu Jul 18th has been shifted to Tue 23rd (all day). OTG0051004 , GocDB.
    • FTS:
      • FTS CMS/LHCb upgrade to v 3.9.1 on Wed 17th Jul from 9:30 to 12:30 CEST (OTG:0051317). The upgrade requires a downtime due to the underlying DB schema upgrade.
      • FTS: VMs reboot due to HVs reboot campaign, Thu 18th Jul (OTG:0051230), Mon 22nd July (OTG:0051232) and Wed 24th July (OTG:0051234)
  • CERN databases: NC
  • GGUS: NTR
  • Monitoring:
    • ATLAS DDM / Job accounting dashboards migrated to the new infrastructure (old ones not available anymore, infrastructure behind will be kept running for some months)
    • CMS Job accounting / Task monitoring dashboards migrated to the new infrastructure (old ones not available anymore, infrastructure behind will be kept running for some months)
    • SLS old meter flow has been stopped, new dashboards available in MONIT
  • MW Officer: NC
  • Networks: NC
  • Security: NTR

AOB:

Edit | Attach | Watch | Print version | History: r19 < r18 < r17 < r16 < r15 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r19 - 2019-07-15 - BorjaGarridoBear
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback